# SkyMask: Attack-agnostic Robust Federated Learning with Fine-grained Learnable Masks

### v1

@@ Abstract
Federated Learning (FL) is becoming a popular paradigm for leveraging distributed data and preserving data privacy. However, due to the distributed characteristic, FL systems are vulnerable to Byzantine attacks that compromised clients attack the global model by uploading malicious model updates. Most existing Byzantine-robust FL systems statistically analyze the weights of whole individual model updates uploaded by clients to defend against Byzantine attacks. With the development of layer-level and parameter-level fine-grained attacks, the attacks’ stealthiness and effectiveness have been significantly improved. Due to unawareness or overreaction, the existing model-level defense methods degrade the training efficiency and model performance. To address this problem, we propose SkyMask, a new attack-agnostic robust FL system that leverages fine-grained learnable masks to identify malicious model updates at the parameter-level. Specifically, the FL server applies parameter-level masks to model updates uploaded by clients and trains the masks over a small clean dataset (i.e., root dataset) to learn the subtle difference between benign and malicious model updates in a high-dimension space. Our extensive experiments involve different models on three public datasets under state-of-the-art (SOTA) attacks, where the results show that SkyMask achieves up to 10% higher testing accuracy compared with SOTA defense strategies and successfully defends against attacks with malicious clients of a high fraction up to 80%. In the meantime, the experimental results demonstrate the scalability of our approach and the weak dependence on the data distribution of the root dataset.
I Introduction
Every user click is an oil drop of the digital era that powers up today’s information industry. Massive volumes of data generated by users and collected by online services (e.g., social networks, online sales, and streaming platforms) are used to train machine learning (ML) models (e.g., recommendation algorithms) that enhance user experience and boost profits. However, with increasingly strict data and privacy policies, such as the GDPR by EU and ATT by Apple, web services tracking and collecting user data for centralized training becomes inappropriate. The emergence of federated learning (FL) has addressed such conflicts between ML training and data privacy [1]. Instead of centralizing training data, FL distributes a model to clients, trains the model locally, and aggregates the local models into a global model without leaking any clients’ local data, which has been applied in many web applications, including online recommendation systems and personal healthcare.
@@ I Introduction
However, due to FL’s distributed design, attackers can easily attack the FL system by compromising client participants and smuggling malicious model updates, known as Byzantine attacks [2, 3, 4, 5, 6, 7]. Based on whether the compromised clients poison their local data or models, Byzantine attacks can be classified into data poisoning and model poisoning attacks. Data poisoning attacks corrupt the global model by injecting contaminated training samples to the local dataset [2, 7]. Model poisoning attacks [3, 5] directly manipulate local model updates to compromise the global model. To tackle these attacks, researchers have explored various defense strategies, most of which leverage coarse-grained model-level statistics, distances, and cosine-similarities to detect outlier model updates (i.e., Multi-Krum [8] and FLTrust [9]) or greedily filter out outlier parameters (i.e., Trim [10] and Bulyan [11]).
@@
Unfortunately, research on fine-grained attacking methodologies is becoming trendy due to their precise balancing between attacking stealthiness and effectiveness. Fine-grained model poisoning attacks [4, 12] leverage the varying sensitivities of different neural network layers and parameters to craft adaptive and stealthy attacks. FIBA [13] injects small pixel-wise perturbations in the frequency domain to conduct data poisoning. Existing model-level defense strategies either fail to prevent such fine-grained attacks or spill over the benign clients, which extensively sacrifices training efficiency and model quality.
@@
This paper proposes SkyMask, a new Byzantine-robust FL system that defends against agnostic attacks with fine-grained learnable masks at the parameter-level. Specifically, after collecting local models from participant clients, the FL server freezes and multiplies each of the local model parameters with a learnable mask variable. Then, all masked local models are trained on a small and clean data set (i.e., the root dataset [9]) together. During the training on the root dataset, the mask variables applied to the frozen local model parameters learn to improve the model accuracy. Particularly, the masks applied to malicious models are trained to correct their potential misbehaviors at the parameter-level. Since utilizing fine-grained learnable masks does not enforce customization for specific attacks or refactoring the FL training objective functions, SkyMask gains three superior capabilities beyond existing defense strategies: 1) effective and efficient detection of fine-grained stealthy attacks, 2) attack-agnostic defense against a wide spectrum of attack methods, and 3) strong compatibility and complementarity with existing FL and defense methodology. Our main contributions are as follows:
@@
  • We design fine-grained and computation-efficiency learnable masks that can capture malicious models’ characteristics in a uniform high-dimension space.
@@
  • We develop SkyMask, a new attack-agnostic Byzantine-robust FL system, which is the first work applying fine-grained learnable masks to detect malicious clients and defend against Byzantine attacks.
@@
  • We empirically evaluate SkyMask on various benchmarks under six SOTA attacking methods and compare them with existing defense methods, where experimental results show that SkyMask achieves up to 10% higher testing accuracy compared with existing defenses.
II Preliminaries & Related Work
II-A Federated Learning
@@
A typical FL system includes $n$ clients and a server. Each client $i$ has a local dataset $D_{i}$, $i=1,\dots,n$. In each communication round $t$, the server selects a subset of the clients $\mathcal{N}_{t}$ to execute the following three steps: 1) Model distribution: The server distributes the global model $W_{t}$ to the selected clients. 2) Local training: The clients receive global model $W_{t}$ as the initial model $W^{i}_{t+1}$ for local training, and repeat $W_{t+1}^{i}:=W^{i}_{t+1}-\beta\nabla f(W^{i}_{t+1};D_{i})$ for $l$ local iterations, where $f(\cdot;\cdot)$ is the empirical loss function and $\beta$ denotes the local learning rate. The corresponding local model update is $\Delta W^{i}_{t+1}=W_{t}-W^{i}_{t+1}$, so for the server, it is equivalent to the local model $W^{i}_{t+1}$. 3) Model aggregation: After the training, each client $i$ uploads its model weights $W_{t+1}^{i}$ to the server for aggregation. FedAvg [1] performs weighted model averaging to update the global model: $W_{t+1}=\sum_{i\in\mathcal{N}_{t}}\frac{|D_{i}|}{\sum_{i\in\mathcal{N}_{t}}|D_{i}|}W_{t+1}^{i}$, where we set $|\mathcal{N}_{t}|=n$ for simplicity.
II-B Threat Model & Byzantine Attacks on FL
@@
Malicious clients compromised by an attacker participate in the FL process and have access to the knowledge of neural network models, learning rates, and objective functions. Besides, the attacker can fully control these clients’ activities (e.g., modifying the model updates) and has complete access to their local datasets [4, 9, 12].
@@
Existing Byzantine attacks on FL can be categorized into untargeted, targeted, or backdoor based on the adversary’s goal: 1) Untargeted attacks seek to corrupt the global model and minimize its accuracy on any test input [2, 4, 12, 5]. For example, Fang [4] adaptively trades off the attacking stealthiness and effectiveness at a fine granularity as an optimization problem. 2) Targeted attacks aim to minimize the accuracy on the specific test inputs, and maintain high accuracy on the rest [3, 14]. 3) Backdoor attacks reduce the utility of the global model on test inputs that contain a specific signal called trigger [15, 6].
II-C Byzantine-robust FL Algorithms
@@
II-C1 Model-level defense strategies
FLDetector [16] determines whether a client is malicious by performing the consistency analysis on its model updates. Specifically, FLDetector predicts a client’s model update for each iteration based on historical records and compares it with the received model update. If they are inconsistent in multiple iterations, the client is probably malicious. FLTrust [9] trains a root model with a small clean root dataset at the server in each round, and measures the cosine similarity between a local model and the root model. The similarity metric dictates the weight in global model aggregation. Tolpegin [2] detects malicious participants by analyzing the characteristics of model updates. It standardizes all model updates, uses the principal component analysis (PCA) for dimension reduction, and identifies malicious clients via clustering algorithms. Krum [8] assumes the number of malicious clients is known as $n_{m}$. The server computes a score $s^{i}_{t+1}$ for each model update from the sum of squares of the distance between this model update and the $n-n_{m}-2$ nearest ones. Then, the one with the smallest score is selected as the global model update for the next round. Multi-Krum (MKrum) [8], a Krum variant, selects $c$ model updates with the lowest scores, and then aggregates them for a new global model.
II-C2 Greedy parameter-level filtering strategies
Trimmed-Mean (Trim) and Median [10] are coordinate-wise aggregation rules assuming the number of malicious clients $n_{m}$ is known, remove the parameters of the smallest and largest $n_{m}$ values, and average the remaining. Median retains the corresponding coordinate’s median value as the new global model’s value. Bulyan [11] also relies on a known $n_{m}$ and requires $n\geq 4n_{m}+3$ to ensure its robustness. It first iteratively applies MKrum to select $\theta$ local models, where $\theta<n-2n_{m}$, and uses a variant of Trim to aggregate the $\theta$ selected local models.
@@ II-C2 Greedy parameter-level filtering strategies
However, existing robust FL strategies can hardly defend against those fine-grained attacks without sacrificing training efficiency and model quality by utilizing coarse-grained similarity checking or greedy parameter-wise filtering, which motivates us to explore malicious client detection with fine-grained learnable masks.
II-D Motivating Learnable Masks
(a) The PCA of model updates
(b) The PCA of masks
Figure 1: Visualizing model updates and masks with PCA.
Binary masks have been widely used in deep learning and demonstrated the capability to capture neural network models’ intrinsic characteristics at the parameter-level. For example, Zhou et al[17] prune deep neural network models with binarized masks to improve training efficiency. FedMask [18] utilizes learnable binary masks to enable each device to learn a personalized and structured sparse neural network model that runs efficiently on devices.
We use a realistic experiment to show how learnable masks can be applied to detect malicious clients. In a swarm of 100 FL clients, we compromise 20 as malicious clients to apply Min-Sum attack. We train a four-layer CNN model on Fashion-MNIST dataset with the non-IID setting described in §IV-A1. We visualize the CNN model updates of all clients in one iteration using PCA in Figure 1(a), which shows how the fine-grained Min-Sum attack hides the malicious model updates among the benign updates. Hence, analyzing local model updates with PCA is insufficient to distinguish benign and malicious clients.
To detect such stealthy attacks, we apply fine-grained learnable masks that capture the parameter-level characteristics of local models. We aggregate all masked local models and train the masks together on a small root dataset, where the learnable masks learn to enhance the benign models and correct the malicious models (details in §III-B). Figure 1(b) shows the visualization of trained masks using PCA, where the masks applied to the malicious model updates are easily distinguished among all masks. This phenomenon reveals the capability of learnable masks that captures the difference between benign and malicious model updates.
II-E Root Dataset
@@
Considering that the server has no root of trust to decide whether a model update is malicious or not, FLTrust [9] was the first to introduce the concept of root dataset. It can take advantage of the root dataset even if the distribution diverges from the overall data distribution or the size is less than one hundred. Therefore, it is easy for the server and FL manager to manually collect or generate a small root dataset. The rationality for using root datasets has been verified in several other approaches [19, 20].
III SkyMask’s Design
III-A Goals
@@
We aim to design a server-based robust FL system. The server has a small representative root dataset following a previous work [9]. Moreover, this root dataset can be imbalanced without a major bias (analyzed in §IV-D1). We allow reasonable additional computation costs and extra storage space to defend against Byzantine attacks. We aim to detect all the malicious model updates and filter them out at each communication round to eliminate the impact of attacks. We should also keep as many benign clients as possible in the filtered FL client set to train a final global model with sufficient local data for satisfactory accuracy. For the proposed Byzantine-robust FL system, it should promise the following three features:
  • Robustness. The system should preserve the global model accuracy when attacked by malicious clients. In particular, we should maintain a high malicious client detection accuracy and a low misidentification rate of benign ones (i.e., false positive rate). The system should also stay effective on different datasets and model structures under different attacks.
  • Efficiency. The system should not introduce extra computation and communications overhead to the clients, which are often resource-constrained devices in FL. Besides, when performing malicious client detection and defense on the server, the system should only incur a proper computation complexity.
  • Scalability. The system should still support a large scale of clients and work under a high fraction of malicious clients (e.g., 80%) since FL systems usually involve a large number of client devices.
III-B SkyMask: Overview
@@
When initializing the FL process, the server creates an initial global model and builds a root dataset as [9]. Figure 2 presents SkyMask’s workflow, where each communication round has six steps: Step 1, the server distributes the global model parameters to each client. Step 2, the clients load the parameters, train the local models and send back these model updates. Step 3, the server freezes received local models and applies learnable masks to each local model update. Step 4, all the masked local models are trained together on the root dataset to converge. Step 5, the server detects and removes the malicious clients by clustering all trained masks. Step 6, the server aggregates the remaining model updates into a new global model.
@@
The key component of SkyMask is the detection mechanism based on learnable masks. Specifically, SkyMask constructs a same-size fine-grained mask for each local model update and trains these masks by optimizing the aggregation result of masked local models until they converge. Byzantine attacks take impairing global model performance as an optimization objective, so the poisoned model updates tend to make the global model far from normal behavior. Therefore, to correct potential misbehaviors of malicious models, the masks applied to the malicious models learn to mitigate the side effects of poisoned parameters through training on the root dataset. In this way, the learnable masks form a high-dimensional representation space to capture the characteristics of both malicious and benign model updates. Then, the clusters formed by masks are the basis for determining whether a client is malicious. After removing all the detected malicious model updates, SkyMask calculates the average of remaining model updates and uses it to update the global model.
III-C The Complete SkyMask Algorithm
@@
Figure 2: SkyMask’s workflow.
III-C1 Mask initialization and training
In the first stage, the server assigns a learnable mask $m_{i}$ of the same size as the model for each client $i$ and initializes them all with $1$s. The model aggregated from masked local models $\tilde{W}_{t+1}$ (called aggregated masked model) is computed by averaging the masked local models $m_{i}\odot W^{i}_{t+1}$. Then, the masks start training on the root dataset with frozen local models. At the beginning of iteration $t+1$, $\tilde{W}_{t+1}$ is calculated as:
$W^{i}_{t+1}=W_{t}-\Delta W^{i}_{t+1},$
@@ III-C1 Mask initialization and training
Intuitively, we know that if we fix all masks as zeros except for one mask to be trained, there is also a solution for this optimization problem. If $\tilde{W}_{t+1}$ converges to the fixed optimal model $\hat{\tilde{W}}_{t+1}$, the problem is equivalent to solving a multivariate primary equation, which has infinite solutions. Moreover, suppose the mask value is not limited. In that case, some masks have very large values, and others are so small for the same dimension that these masks can not effectively characterize whether the local model is benign. So we cannot detect malicious clients with the help of this type of mask, and we should redesign the masks to show the contribution of each local model update in the optimized model. Binary can directly determine whether a parameter of model update participates in the global model, and binary masks have shown effectiveness in model pruning. So we use binary masks to extract the parameter-level characteristics of local models. If the malicious model parameters attempt to corrupt the aggregated model, the optimization process prefers to reject them and sets the corresponding parameters of the mask as 00.
III-C2 Approximation of binary masks
Binary parameters cannot be easily optimized by a gradient descent algorithm. Thus, we approximate the binary masks in training by setting the mask with real values and applying a sigmoid function $\sigma(\cdot)$ instead of a hard threshold to reduce the enormous gradient variance. Furthermore, it is easy for the gradient to be back-propagated to the real-value masks. The approximation of a binary mask is calculated as $\tilde{m}_{i}=\sigma(m_{i})$. Since the mask’s limitation is changed, the original Equation 1 is unsuitable. To adapt for the approximated binary masks, $\tilde{W}_{t+1}$ is re-written as:
@@ III-C2 Approximation of binary masks
The mask training task is not different from the main task in FL except for the variables. By using the stochastic gradient descent algorithm (SGD) to optimize these masks, the objective function is:
@@
During training the masks, masks are updated in each mask training iteration as follows:
@@
where $\gamma$ represents the mask learning rate.
This process will not stop until these masks converge. After the real-value masks converge, the final results of binary mask $\hat{m}_{i}$ are computed as Equation 5.
@@
III-C3 Mask clustering and classification
These binary masks can sense the fine-grained poisoning attack and represent the parameter-level characteristics of each client’s model update. We use the PCA algorithm to reduce the dimensionality to $d$. Since the maximum dimensionality that the masks can be reduced to is limited by the number of masks $n$ and the dimensionality of masks $V$, we set $d=2$, which is the minimum possible number of clients in the FL system. Then, we apply the Gaussian mixture model for clustering and classification. If there are no attacks, the clustering result is only one cluster, and the server aggregates all the benign local models.
If the fraction of malicious clients is lower than 50%, the system can treat the larger cluster in two clusters as the benign client set. To cope with more general situations, SkyMask introduces a trusted root model $W^{r}_{t+1}$, which is initialized by $W_{t}$ and “locally” trained $l$ iterations on the root dataset as other clients do:
@@ III-C3 Mask clustering and classification
Then the server lets it join in the model set before the mask training stage, assigns one mask for it and trains this mask together with other masks. When all the masks converge, the clients represented by the masks in the same cluster as the mask corresponding to the trusted root model are considered benign. This way, SkyMask can handle various malicious fractions from zero to very high. In order not to interfere with the global model convergence, we should emphasize that the trusted root model is only used for malicious model detection and will not be included in the benign model aggregation phase.
SkyMask plays $R$ rounds for server-clients communication, and each round contains six steps, as §III-B describes. The critical point of SkyMask is our fine-grained malicious client detection strategy mentioned above, which can provide the remaining benign client set $U_{b}$ for aggregating a new global model. The global model for the next communication round $W_{t+1}$ is calculated as:
@@
where $\alpha$ denotes the global learning rate. It should be noted that SkyMask is more concerned with malicious client detection and can be added as an additional module to the aggregation algorithm. Therefore, it can even enhance the existing Byzantine-robust aggregation algorithms. If it completes the malicious detection task, removing only the malicious model updates and leaving the benign model updates untamed, its convergence will be consistent with the original aggregation algorithm.

### v2

@@ Abstract
Federated Learning (FL) is becoming a popular paradigm for leveraging distributed data and preserving data privacy. However, due to the distributed characteristic, FL systems are vulnerable to Byzantine attacks that compromised clients attack the global model by uploading malicious model updates. With the development of layer-level and parameter-level fine-grained attacks, the attacks’ stealthiness and effectiveness have been significantly improved. The existing defense mechanisms solely analyze the model-level statistics of individual model updates uploaded by clients to mitigate Byzantine attacks, which are ineffective against fine-grained attacks due to unawareness or overreaction. To address this problem, we propose SkyMask, a new attack-agnostic robust FL system that firstly leverages fine-grained learnable masks to identify malicious model updates at the parameter level. Specifically, the FL server freezes and multiplies the model updates uploaded by clients with the parameter-level masks, and trains the masks over a small clean dataset (i.e., root dataset) to learn the subtle difference between benign and malicious model updates in a high-dimension space. Our extensive experiments involve different models on three public datasets under state-of-the-art (SOTA) attacks, where the results show that SkyMask achieves up to 14% higher testing accuracy compared with SOTA defense strategies under the same attacks and successfully defends against attacks with malicious clients of a high fraction up to 80%. Code is available at https://github.com/KoalaYan/SkyMask.
1 Introduction
Federated learning (FL) [22] is a distributed machine learning paradigm that addresses the conflicts between ML training and data privacy. Instead of centralizing training data, FL distributes a global model to clients, trains the model locally, and aggregates the local models into a new global model without leaking any clients’ local data, which has been applied in many computer vision applications, including large-scale visual classification [11], object detection [20], medical imaging [10, 18], and many others [35, 19, 17].
@@ 1 Introduction
However, due to FL’s distributed design, attackers can easily attack the FL system by compromising client participants and smuggling malicious model updates, known as Byzantine attacks [3, 2, 32, 7, 28, 26]. To tackle Byzantine attacks, researchers have explored various defense strategies, most of which leverage coarse-grained model-level statistics to detect outlier model updates [4, 5, 28, 27, 24] or greedily filter out outlier parameters [33, 8].
@@
Unfortunately, fine-grained attacking methodologies utilize their precise balancing between attacking stealthiness and effectiveness to circumvent existing defense mechanisms, e.g., Fang attack [7] and AGR-agnostic attack [28] leverage the varying sensitivities of different neural network layers and parameters to craft adaptive and stealthy attacks. Existing model-level defense strategies either fail to prevent such fine-grained attacks or spill over the benign clients, which extensively sacrifices training efficiency and model quality. We conduct a real-world experiment to show the poor performance of the model-level defense [30] under a fine-grained attack [28] (detailed settings provided in supplementary §A.5). Existing defense method exhibits 18% testing accuracy degradation. We visualize model updates of all clients in a single iteration by the principal component analysis (PCA) as [30]. As depicted in Fig. LABEL:fig:toy_model, the fine-grained attack camouflages the malicious updates within benign ones, underscoring the insufficiency of analyzing local model updates solely at the model level.
To explore whether binary masks [37, 14] can effectively detect malicious clients, we simultaneously train the binary masks for each model update on the server in the aforementioned toy experiment. Fig. LABEL:fig:toy_mask illustrates the visualization of trained masks using PCA, where masks applied to malicious model updates distinctly stand out among all masks. This highlights the capability of learnable masks to differentiate between benign and malicious model updates, and this method improves the testing accuracy by 18%.
@@
Based on this observation, we propose SkyMask, a new Byzantine-robust FL system that defends against agnostic attacks with fine-grained learnable masks at the parameter level. Specifically, after collecting local models from participant clients, the FL server freezes and multiplies each of the local model parameters with a learnable mask variable. Then, all masked local models are trained on a small and clean dataset (i.e., the root dataset [5]) together. During the training process on the root dataset, the mask variables, which are applied to the frozen local model parameters, learn to improve the model accuracy. Particularly, the masks applied to malicious models are trained to correct their potential misbehaviors at the parameter level. Since utilizing fine-grained learnable masks does not enforce customization for specific attacks or refactoring the FL training objective functions, SkyMask gains three superior capabilities beyond existing defense strategies: 1) effective and efficient detection of fine-grained stealthy attacks, 2) attack-agnostic defense against a wide spectrum of attack methods, and 3) strong compatibility and complementarity with existing FL and defense methodology. To sum up, our main contributions are as follows:
@@
  • We propose fine-grained learnable masks that can capture models’ intrinsic characteristics at the parameter level in a uniform high-dimension space.
@@
  • We design SkyMask, a new attack-agnostic Byzantine-robust FL system, which applies fine-grained learnable masks to detect malicious clients and defend against Byzantine attacks.
@@
  • We empirically evaluate our SkyMask on various benchmarks under seven SOTA attacking methods and compare them with existing defense methods, where experimental results show that SkyMask achieves up to 14% higher testing accuracy compared with existing defenses.
2 Preliminaries & Related Work
2.1 Federated Learning
@@
A typical FL system includes $n$ clients and a server. Each client $i$ has a local dataset $D_{i}$, $i=1,\dots,n$. In each communication round $t$, the server selects a subset of the clients $\mathcal{N}_{t}$ to execute the following steps: 1) Model distribution: The server distributes the global model $W_{t}$ to the selected clients. 2) Local training: The clients receive global model $W_{t}$ as the initial model $W^{i}_{t+1}$ for local training, and repeat $W_{t+1}^{i}:=W^{i}_{t+1}-\beta\nabla f(W^{i}_{t+1};D_{i})$ for $l$ local iterations, where $f(\cdot;\cdot)$ is the empirical loss function and $\beta$ denotes the local learning rate. The corresponding local model update is $\Delta W^{i}_{t+1}=W_{t}-W^{i}_{t+1}$. 3) Model aggregation: After the training, each client $i$ uploads its model weights $W_{t+1}^{i}$ to the server for aggregation. FedAvg [22] performs weighted model averaging to update the global model: $W_{t+1}=\sum_{i\in\mathcal{N}_{t}}\frac{|D_{i}|}{\sum_{i\in\mathcal{N}_{t}}|D_{i}|}W_{t+1}^{i}$, where we set $|\mathcal{N}_{t}|=n$ for simplicity.
2.2 Threat Model & Byzantine Attacks on FL
@@
Malicious clients compromised by the attacker participate in the FL process and have access to the knowledge of neural network models, learning rates, and objective functions. Besides, the attacker can fully control these clients’ activities (e.g., modifying the model updates) and has complete access to their local datasets [7, 5, 28].
@@
Existing Byzantine attacks on FL can be categorized into untargeted and targeted based on the adversary’s goal: 1) Untargeted attacks seek to corrupt the global model and minimize its accuracy on any test input [30, 7, 28, 2]. Specifically, fine-grained attacks, represented by Fang attack [7] and AGR-agnostic attack [28], adaptively trade off the attacking stealthiness and effectiveness at a fine granularity as an optimization problem, which preserves the effects of poisoning with as few model-level outliers as possible. 2) Targeted attacks aim to reduce the utility of the global model on attacker-specified tasks [3, 29, 1, 32, 15, 21, 25, 34].
2.3 Byzantine-robust FL Algorithms
@@
Model-level defense strategies. Some defense methods take model-level distance as a basis for determining whether a client is malicious, e.g., Euclidean distance in Krum [4], or cosine distance in DeepSight [27] and FLAME [24]. FLTrust [5] trains a root model with a small root dataset, and takes the cosine similarity between a local model and the root model as the weight in aggregation. Tolpegin defense [30] identifies malicious model updates by employing PCA dimension reduction and clustering. FLDetector [36] predicts a client’s model update based on historical data and compares it with the received update. It employs consistency analysis on model updates to identify malicious clients.
Greedy parameter-level filtering strategies. Trimmed-Mean (Trim) [33] is a coordinate-wise aggregation rule assuming the number of malicious clients $n_{m}$ is known, removes the parameters of the smallest and largest $n_{m}$ values, and averages the remaining.
@@
However, existing robust FL strategies can hardly defend against those fine-grained attacks without sacrificing training efficiency and model quality by utilizing coarse-grained detection or greedy parameter-wise filtering, which motivates us to explore malicious client detection with fine-grained learnable masks.
2.4 Root Dataset
@@
Considering that the server has no root of trust to decide whether a model update is malicious or not, FLTrust [5] is the first to introduce the concept of root dataset. It can take advantage of the root dataset even if the distribution diverges from the overall data distribution or the size is less than one hundred, which is easy for the server and FL manager to manually collect or generate a small root dataset. The rationale for utilizing root datasets has been validated in numerous other studies [9, 23, 31, 16]. We evaluate the impact of the root dataset’s data distribution in supplementary §C.1.
3 Methodology
@@
We aim to design a server-based robust FL system. The server has a small representative root dataset following previous work [5]. Moreover, this root dataset can be imbalanced. For the proposed Byzantine-robust FL system, it should promise the following three features: (1) Robustness. The system should preserve the global model accuracy under attacks. In particular, we should maintain a high malicious client detection accuracy and a low misidentification rate of benign ones to eliminate the impact of attacks. (2) Generality. The system should also stay effective on different datasets and model structures under different attacks. (3) Efficiency. The defense system is designed on the server instead of resource-constrained client. Besides, when performing malicious client detection and defense on the server, we allow reasonable additional computation costs and extra storage space.
3.1 Overview
Figure 4: SkyMask’s workflow.
@@ 3.1 Overview
When initializing the FL process, the server creates an initial global model and builds a root dataset as [5]. Fig. 4 presents SkyMask’s workflow, where each communication round has six steps: Step 1, the server distributes the global model parameters to each client; Step 2, the clients load the parameters, train the local models, and send back these model updates; Step 3, the server freezes and multiplies each local model parameter with a learnable mask variable; Step 4, all the masked local models are trained together on the root dataset to converge; Step 5, the server detects and removes the malicious clients by clustering all trained masks; Step 6, the server aggregates the remaining model updates into a new global model.
@@
The key component of SkyMask is the learnable mask-driven detection. Specifically, SkyMask constructs a same-size fine-grained mask for each local model update and trains these masks by optimizing the aggregation result of masked local models until they converge. Byzantine attacks take impairing global model performance as an optimization objective, so the poisoned model updates tend to make the global model far from normal behavior. Therefore, to correct potential misbehaviors of malicious models, the masks applied to the malicious models learn to mitigate the side effects of poisoned parameters through training on the root dataset. In this way, the learnable masks form a high-dimensional representation space to capture the characteristics of both malicious and benign model updates. Then, the clusters formed by masks are the basis for determining whether a client is malicious. After removing all the detected malicious model updates, SkyMask calculates the average of remaining model updates and uses it to update the global model.
3.2 SkyMask Algorithm
@@
3.2.1 Mask initialization and training.
In the first stage, the server freezes all the model updates and assigns a learnable mask $m_{i}$ of the same size as the model for each client $i$ and initializes them all with $1$s. The model aggregated from masked local models $\tilde{W}_{t+1}$ (called aggregated masked model) is computed by averaging the masked local models $m_{i}\odot W^{i}_{t+1}$. Then, the masks start training on the root dataset with frozen local models. At the beginning of iteration $t+1$, $\tilde{W}_{t+1}$ is calculated as:
@@
If there is no constraint on the value range of mask variables during the mask training, some masks can have very large values, while others can be so small for the same dimension when these masks converge. Huge differences in magnitude between masks can affect the detection of malicious clients. Therefore, we redesign the masks to show the contribution of each local model update in the optimized model.
We use binary masks to extract the parameter-level characteristics of local models because binary can directly determine whether a parameter of model update participates in the global model. Since binary parameters are not derivable, we approximate the binary masks in training by setting the mask with real values and applying a sigmoid function $\sigma(\cdot)$ instead of a hard threshold to reduce the enormous gradient variance [14]. Furthermore, it is easy for the gradient to be back-propagated to the real-value masks. The approximation of a binary mask is calculated as $\tilde{m}_{i}=\sigma(m_{i})$, which is ranged $(0,1)$.
Since the mask’s limitation is changed, the original Equation 1 is unsuitable. To adapt for the approximated binary masks, $\tilde{W}_{t+1}$ is re-written as:
@@
The mask training task is not different from the main task in FL except for the variables. The objective function is:
@@
The masks are updated in each mask training iteration as follows:
@@
where $\gamma$ represents the mask learning rate. This process will not stop until these masks converge. After the real-value masks converge, the final results of binary mask $\hat{m}_{i}$ are computed as:
@@
3.2.2 Mask clustering and classification.
These binary masks can sense potential attacks at a fine granularity and represent the parameter-level characteristics of each client’s model update. Then, we apply the Gaussian mixture model for clustering and classification. If there are no attacks, the clustering result is only one cluster, and the server aggregates all the benign local models.
To determine which cluster is benign, SkyMask introduces a trusted root model $W^{r}_{t+1}$, initialized by $W_{t}$ and “locally” trained $l$ iterations on the root dataset as other clients:
@@
The server lets it join in the model set before the mask training stage, assigns one mask for it and trains this mask together with other masks. When all the masks converge, the clients represented by the masks in the same cluster as the mask corresponding to the trusted root model are considered benign. In this way, SkyMask can handle various malicious fractions from zero to very high. In order not to interfere with the global model convergence, we should emphasize that the trusted root model is only used for malicious model detection and is not included in the benign model aggregation phase.
SkyMask plays $R$ rounds for server-clients communication, and each round contains six steps, as §3.1 describes. The critical point of SkyMask is our fine-grained malicious client detection strategy mentioned above, which can provide the remaining benign client set $U_{b}$ for aggregating a new global model. Given the aggregation algorithm specified as FedAvg, the global model for the next communication round $W_{t+1}$ is calculated as:
@@
where $\alpha$ denotes the global learning rate. If it completes the malicious detection task, removing only the malicious model updates and leaving the benign model updates untamed, its convergence will be consistent with the original aggregation algorithm. SkyMask serves as an efficient method for identifying malicious clients within an FL framework. Its modular design allows for integration as an adjunct component into diverse aggregation algorithms, ensuring compatibility across a spectrum of existing aggregation techniques.
3.3 Why SkyMask Works